1. An object has a mass of 2kg is dropped from the top of a building 30 meters tall. the initial velocity is zero. As it falls, the object encounters air resistance that is equal to 1/3 v. what is the velocity and attitude of the object after 1.5 seconds? Ans. 13.02m/s; 19.83m
  2. A tank initially contains 50 gal of fresh water. At t=0, brine solution containing 2 lb of salt per gallon is poured into the tank at the rate of 12 gal/min, while the well stirred mixture leaves the tank at the rate of 8 gal/min. what is the amount of salt at the end of 5 minutes? How long will it take to obtain an amount of 50 lb? ans. 88098 lb; 2.47 min
1
Expert's answer
2021-11-17T15:39:54-0500

1. The second Newton's Law


"mv'=mg-\dfrac{1}{3}v"

Given "m=2\ kg, g=9.81\ m/s^2"


"v'+\dfrac{1}{6}v=9.81"

"v'=-\dfrac{1}{6}(v-58.86)"

"\dfrac{dv}{v-58.86}=-\dfrac{1}{6}dt"

integrate


"v(t)=58.86+c_1e^{-t\/6}"

"v(0)=0=>c_1=-58.86"

"v(t)=58.86-58.86e^{-t\/6}"

"v(1.5)=58.86-58.86e^{-1.5\/6}"

"v(1.5)=13.02\ m/s"

"v(t)=-h'(t)"

"h(t)=-\int v dt=-\int(58.86-58.86e^{-t/6})dt"

"=-58.86t-353.16e^{-t\/6}+c_2"

"h(0)=30=-353.16+c_2=>c_2=383.16"

"h(t)=-58.86t-353.16e^{-t\/6}+383.16"

"h(1.5)=-58.86(1.5)-353.16e^{-1.5\/6}+383.16"

"h(1.5)=19.83\ m"

2. Let "s(t) =" amount, in lb of salt at time "t." Then we have

"\dfrac{ds}{dt}="(rate of salt into tank) − (rate of salt out of tank)


"\dfrac{ds}{dt}=2\cdot12-\dfrac{8s}{50+(12-8)t}"

So we get the differential equation


"\dfrac{ds}{dt}=24-\dfrac{8s}{50+4t}"

"\dfrac{ds}{dt}+\dfrac{4s}{25+2t}=24"

Integrating factor


"\mu(t)=e^{\int 4dt/(25+2t)}=(25+2t)^{2}"

"\dfrac{d}{dt}((25+2t)^{2}s)=24(25+2t)^{2}"

"\int d((25+2t)^{2}s)=24\int(25+2t)^{2}dt"

"(25+2t)^{2}s=24(\dfrac{1}{6})(25+2t)^{3}+c_1"

"s(t)=4(25+2t)+c_1(25+2t)^{-2}"

"s(0)=100+c_1(25)^{-2}=0"

"c_1=-62500"

"s(t)=4(25+2t)-62500(25+2t)^{-2}"

"s(5)=4(25+2(5))-62500(25+2(5))^{-2}"

"s(5)=88.98\ lb""s(t)=4(25+2t)-62500(25+2t)^{-2}=50"

Let "y(t)=s(t)-50"

"y(t)=4(25+2t)-62500(25+2t)^{-2}-50"

Solve graphically the equation "y=0"

"t=2.47\ min"